OpenStudy (anonymous):

Find an equation of the tangent line to the given curve at the specified point y= (2x)/(x+1) at (0,0)

OpenStudy (anonymous):

I meant (1,1) sorry

OpenStudy (anonymous):

first step would be to find the derivative of y:  which gives:\[\frac{ 2 }{ (x+1)^{2} }\] At (1,1),gradient is \[\frac{ 1 }{ 2 }\] therefore,we have: y=mx + c 1=1(1/2) + c c= 1/2 equation of tangent: y= (1/2)x + (1/2)